OpenStudy (anonymous):

Determine the coordinates of the x-intercepts for the quadratic equation -x2 + 5x - 4 = 0

OpenStudy (anonymous):

x2-5x+4=0 multiply b.s. by -1 x2-4x-x+4=0 x(x-4)-1(x-4)=0 x=1 and x=4

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np :)